# Is In-Context Learning Sufficient for Instruction Following in LLMs?

### v2

@@ 1 Introduction
Analysis of in-context alignment. In this work, we extend the evaluation of the Urial prompt strategy proposed by Lin et al. (2024) across several base models, including GPT-4-Base [^1] We received access to the base GPT-4 model via the OpenAI Researcher Access Program. and on established instruction following benchmark MT-Bench (Zheng et al., 2023), see Table 1. First, we show that, although Urial achieves reasonable performance, it still lags behind instruction-fine-tuned models. Next, to better understand the success but also weaknesses of in-context alignment, we analyze which are the most relevant ingredients in Urial. We find that the decoding parameters in the LLMs generation (temperature, sampling scheme, etc.) may crucially influence the performance, and the optimal configuration allows even base models to achieve good scores on MT-Bench. Moreover, we show that randomly sampled triplets of examples from the highly curated instruction dataset SkillMix (Kaur et al., 2024) can achieve similar, or even better, results than Urial when used for in-context alignment.
@@
Many-shot in-context alignment. Next, we test various strategies to improve in-context alignment, leveraging recent models with extensive context windows which allow for longer in-context prompts. In particular, we study the effect of many-shot in-context learning by adding high-quality demonstrations from existing instruction datasets. Unlike what suggested by Lin et al. (2024), this approach can improve upon Urial, but only when using high-quality examples, as those from SkillMix. However, it is still not sufficient to fully close the gap with aligned LLMs, as the performance plateaus after 10-30 in-context examples. This behavior is in contrast to many-shot ICL for tasks like summarization (Narayan et al., 2018), translation (Costa-jussà et al., 2022), or classification (Li et al., 2024), where providing many examples is clearly beneficial (Agarwal et al., 2024; Bertsch et al., 2024). We further test a simple greedy algorithm to select the in-context examples which optimize the MT-Bench score. This selection scheme outperforms, with 1 to 3 additional demonstrations, the many-shot approach with random samples, and allows to further reduce the distance from in-context aligned models to fine-tuned models.
@@
ICL vs IFT for alignment. While our experiments suggest that ICL cannot match the instruction-following performance of models aligned through costly fine-tuning, possibly using preference data, it remains an open question whether ICL can compete with or outperform IFT in the low-data regime. We provide an extensive evaluation on multiple LLMs and datasets of both approaches when varying the question-answer dataset size between 3 and 4000 examples. We observe that, with high-quality data, ICL and IFT achieve almost identical 1st-turn MT-Bench score. Surprisingly, in the 2nd-turn score, IFT clearly outperforms ICL, with ICL performing even worse than the base model. This suggests that ICL suffers from some heavy overfitting to the style of the examples shown in context.
@@
Overall, these results, complemented by several ablation studies, provide a more nuanced picture of ICL as an alignment technique compared to previous works. Moreover, our comparison of ICL to IFT when using the same data bridges a gap in the literature about understanding these orthogonal approaches for adapting pre-trained LLMs into conversational models.
@@ 2.1 Systematic evaluation of Urial
Lin et al. (2024) propose to use a prompt with as few as three constant stylistic examples and a carefully designed set of rules to align base LLMs with ICL. The highly curated examples always begin with affirming the user query, then proceed to answer the query by enumerating necessary items and steps, and conclude with an engaging summary. The set of rules first elucidate the scenario and format of the subsequent interaction between the human user and the AI assistant, then outline multiple qualitative aspects, including helpfulness, honesty, and harmlessness, which the base LLMs should adhere to when answering queries.
We compare the performance of several base models with the Urial in-context prompt to that of their instruction fine-tuned versions. Table 1 shows the results on MT-Bench (Zheng et al., 2023), one of the most popular benchmarks for instruction following ability of LLMs. We note that Lin et al. (2024) originally tested Urial on their own dataset, in which only $8\%$ of the examples are obtained from MT-Bench. We compare several LLMs of different sizes and capabilities, ranging from Llama-2-7B (Touvron et al., 2023) to the base GPT-4 model (OpenAI, 2023). For simplicity, we consider the official instruction-tuned LLMs, i.e., provided from the same source of the base models, even if there exist third-party fine-tuned versions which can achieve higher scores. We observe that base models with Urial achieve competitive performance on the 1st-turn score, but cannot match that of their instruction fine-tuned counterparts in all cases except for Llama-2-70B and Mistral-7B-v0.1 (both originally included in Lin et al. (2024), unlike most others). Conversely, the 2nd-turn score with Urial is significantly worse than for instruction-tuned LLMs: we hypothesize that this is because no multi-turn demonstrations are given in context. Indeed, Zhou et al. (2023) show that, even in IFT, having a few multi-turn training examples significantly improve the performance of the model on multi-turn conversations. Because of this fact, in the rest of the paper, we mainly focus on the 1st-turn score and single-round conversations, and track of 2nd-turn performance for completeness. Finally, we detail the breakdown of the results over categories in Table 7 (see App. C.3 for details).
@@ 2.2 What matters for in-context alignment via Urial?
Figure 1: Effect of decoding parameters on the 1st-turn MT-Bench scores across models. We vary temperature and top-$p$, with a fixed repetition penalty of 1.15 as used in the Urial codebase. The heatmaps show that the answering quality of the base model (Mistral-7B-v0.2) with and without Urial is sensitive to the decoding schemes. Conversely, the performance of the instruct model is robust to varying the decoding parameters. Surprisingly, with proper decoding parameters, the base model alone is already capable of following instructions. See the complete results in App. C.2.
Decoding parameters. When computing the performance of Urial, Lin et al. (2024) use decoding parameters (temperature $=0$, top-$p=1$, repetition penalty $=1.15$) which differ from the default ones in MT-Bench (where temperature depends on the topic of the question, top-$p=1$, repetition penalty $=1$). To clarify the influence of these parameters that were not explicitly discussed in Lin et al. (2024), we compute the performance of base LLM, base LLM with Urial, and fine-tuned LLM when varying decoding configurations. In Fig. 1 we fix the repetition penalty to $1.15$, and vary temperature and top-$p$, for the Mistral-7B-v0.2 models. Surprisingly, we notice that with the decoding parameters of Urial (temperature $=0$, top-$p=1$) even the base model without any in-context example achieves reasonable MT-Bench score (6.61 vs 7.00 of Urial). Moreover, the temperature value appears to have the most influence of the performance of the base model. With Urial, almost all configurations provide very similar results, with the exception of temperature $=1$, top-$p=1$ (possibly because it allows the sampling of generated tokens from the tail of the token distribution, thus producing low-quality text). Finally, the aligned model performs similarly across all decoding schemes, with slightly better results than Urial. These results suggest that (1) the decoding parameters are an overlooked factor contributing to the success of in-context alignment, and (2) fine-tuning adjusts the sampling distribution of language models so that even with high-variance decoding configurations the generated text preserves high quality. We provide additional analysis for no repetition penalty (i.e. $=1$) in Fig. 8 and Llama-3.1-8B in Fig. 9 in App. C.2: these extensive experiments further validate, with different base LLMs and settings, that the decoding scheme is a crucial factor for the instruction-following behavior of base models, and also affects, but to lower degree, in-context alignment. Instead, the fine-tuned models are robust to variations in the decoding parameters. Then, for the rest of the experiments we fix the decoding scheme of Urial.
@@ 3.1 Scaling up the number of in-context demonstrations
Finally, we notice that Llama-3.1-8B using three examples from SkillMix attains, on average, better performance than using the three examples of Urial. Overall, there is no significant difference in the scaling behavior between using or not using the Urial demonstrations. Thus, we conclude that the success of in-context alignment is only partially dependent on the demonstrations themselves, provided they are of sufficient quality (see also Sec. 4).
@@ 3.2 Greedy search for effective demonstrations
We add the results of the greedy search to the plots in Fig. 3, computing the true MT-Bench (i.e., with GPT-4 as judge). For both base models, the 4th example found by greedy search is sufficient to match the best 1st-turn score achieved by scaling the in-context examples with random demonstrations (see Sec. 3.1). For Llama-3.1-8B, the 5th and 6th demonstrations further improve the score, while they are not helpful for Mistral-7B-v0.2. In Table 3, we provide the details of these results: the 4th ICL example yields a significant improvement over Urial, e.g., from 6.95 to 7.61 for Llama-3.1-8B, with only a further +0.20 given by the 5th and 6th. Overall, this evaluation further indicates that the improvement in instruction following one can achieve with in-context alignment quickly saturates when increasing the number of examples. In Fig. 4, we show the distribution of the scores (with GPT-4-Turbo as judge) at step of the greedy search when using Llama-3.1-8B as base model (the analog for Mistral-7B-v0.2 in App. C.1). Most demonstrations improve the score when added as the 4th example on top of Urial (as shown in the first plot), but no further progress is observed with additional demonstrations.
@@ 5 Conclusions
In this work, we have first illustrated that ICL via Urial is a good baseline for instruction-following alignment, but with a few limitations: it typically performs slightly worse than IFT on single-turn conversations and does not generalize well to multi-round ones. Next, we have uncovered the key components for IC alignment: e.g. the decoding parameters alone can, surprisingly, make base models reasonably good at instruction following. Adding in-context high-quality demonstrations improves performance beyond what previously suggested (Lin et al., 2024), but not enough to reach LLMs aligned with sophisticated methods (e.g., RLHF). We conjecture that via ICL, the LLM can learn to infer the response style, but the overall learning signal is not sufficiently strong to benefit from a large amount of examples, despite the long context windows of recent LLMs.
@@ Acknowledgements
We are grateful to OpenAI for providing us API credits and access to GPT-4-Base as part of the OpenAI Researcher Access Program. M.A. was supported by the Google Fellowship and Open Phil AI Fellowship. This work was supported by the Swiss National Science Foundation (grant number 212111) and by an unrestricted gift from Google.
@@ A.1 Datasets and base models
We select instruction-following demonstrations for the base LLMs to learn in context from open-sourced instruction fine-tuning datasets: (a) SkillMix-4k (Kaur et al., 2024) is an automated approach for synthesizing diverse, high-quality IFT data. It primarily involves two stages: first leveraging the LLMs to propose a set of critical ”skills” for instruction-following, from which a pair of skills are randomly chosen to facilitate synthetic data generation based on powerful LLMs. (b) Evol-Instruct-70k (Xu et al., 2024) contains 70k training examples with varying complexity and is well-known for the use to build the series of WizardLM models. SkillMix-4k is the primary dataset for all experiments in the paper and Evol-Instruct-70k is solely used for scaling experiments and ablation study of data quality.

### v3

@@ 1 Introduction
Analysis of in-context alignment. In this work, we extend the evaluation of the Urial (the abbreviation of Untuned LLMs with Restyled In-context ALignment) prompt strategy proposed by Lin et al. (2024) across several base models, including GPT-4-Base, [^1] We received access to the base GPT-4 model via the OpenAI Researcher Access Program. and on established instruction following benchmark MT-Bench (Zheng et al., 2023), see Table 1. First, we show that, although Urial achieves reasonable performance, it still lags behind instruction-fine-tuned models. Then, to better understand the success but also weaknesses of in-context alignment, we analyze which are the most relevant ingredients in Urial. We find that the decoding parameters in the LLMs generation (temperature, sampling scheme, etc.) may crucially influence the performance, and the optimal configuration allows even base models to achieve good scores on MT-Bench. Moreover, we show that randomly sampled triplets of examples from the highly curated instruction dataset SkillMix (Kaur et al., 2024) can achieve similar, or even better, results than Urial when used for in-context alignment.
@@
Many-shot in-context alignment. Then, we test various strategies to improve in-context alignment, leveraging recent models with extensive context windows which allow for longer in-context prompts. In particular, we study the effect of many-shot in-context learning by adding high-quality demonstrations from existing instruction datasets. Unlike what suggested by Lin et al. (2024), this approach can improve upon Urial, but only when using high-quality examples, as those from SkillMix. However, it is still not sufficient to fully close the gap with aligned LLMs, as the performance plateaus after 10-30 in-context examples. This behavior is in contrast to many-shot ICL for tasks like summarization (Narayan et al., 2018), translation (Costa-jussà et al., 2022), or classification (Li et al., 2024), where providing many examples is clearly beneficial (Agarwal et al., 2024; Bertsch et al., 2024). We further test a simple greedy algorithm to select the in-context examples which optimize the MT-Bench score. This selection scheme outperforms, with 1 to 3 additional demonstrations, the many-shot approach with random samples, and allows to further reduce the distance from in-context aligned models to fine-tuned models.
@@
ICL vs IFT for alignment. While our experiments suggest that ICL cannot match the instruction-following performance of models aligned through costly fine-tuning, possibly using preference data, it remains an open question whether ICL can compete with or outperform IFT in the low-data regime. We provide an extensive evaluation on multiple LLMs and datasets of both approaches when varying the question-answer dataset size between 3 and 4000 examples. We observe that, with high-quality data, ICL and IFT achieve almost identical 1st-turn MT-Bench score. Surprisingly, in the 2nd-turn score, IFT clearly outperforms ICL, with ICL performing even worse than the base model. This suggests that ICL overfits to the style of the examples shown in context.
@@
Overall, these results, complemented by several ablation studies, provide a more nuanced picture of ICL as an alignment technique compared to previous works. Moreover, our comparison of ICL to IFT when using the same data bridges a gap in the literature about understanding these orthogonal approaches for adapting pre-trained LLMs into conversational models. In summary, our contributions are as follows,
  • Analysis of In-Context Alignment (Sec. 2): We systematically evaluate Urial on a broader set of base LLMs, including GPT-4-Base, with established instruction-following benchmark. Our results indicate that in-context alignment with Urial still underperforms instruction fine-tuning and more sophisticated alignment methods, and a proper decoding scheme is the crucial ingredient behind the empirical success of Urial.
  • Scaling In-Context Alignment (Sec. 3): We find that many-shot ICL with high-quality examples and a simple greedy algorithm can reduce the gap between in-context aligned models to fine-tuned models.
  • First systematic comparison of ICL vs IFT for instruction following (Sec. 4): We show that ICL and IFT with the same number of examples are roughly equivalent for single-turn conversations in the low-data regime. However, IFT generalizes substantially better than ICL when more examples are present, especially for multi-turn conversations.
@@ 2.1 Systematic evaluation of Urial
Background on Urial. Lin et al. (2024) propose to use a prompt with as few as three constant stylistic examples, which follow a curated answering structure that human users are familiar with, and a carefully designed set of rules, highlighting the format and principles AI assistant needs to follow strictly when generating answers, to align base LLMs with ICL. We present the complete Urial prompt in Fig. 9. The highly curated examples always begin with affirming the user query by saying something like “Hi, I’m happy to help you.”, then proceed to answer the query by enumerating necessary items and steps, and conclude with an engaging summary. The set of rules first elucidate the scenario and format of the subsequent interaction between the human user and the AI assistant, then outline multiple qualitative aspects, including helpfulness, honesty, and harmlessness, which the base LLMs should adhere to when answering queries.
Experimental setup. We compare the performance of several base models with the Urial in-context prompt to that of their instruction fine-tuned versions. Table 1 shows the results on MT-Bench (Zheng et al., 2023), one of the most popular benchmarks for instruction following ability of LLMs. We note that Lin et al. (2024) originally tested Urial on their own dataset, in which only $8\%$ of the examples are obtained from MT-Bench. We compare several LLMs of different sizes and capabilities, ranging from Llama-2-7B (Touvron et al., 2023) to the base GPT-4 model (OpenAI, 2023). For simplicity, we consider the official instruction-tuned LLMs, i.e., provided from the same source of the base models, even if there exist third-party fine-tuned versions which can achieve higher scores.
Results. In Table 1 we observe that base models with Urial achieve competitive performance on the 1st-turn score, but cannot match that of their instruction fine-tuned counterparts in all cases except for Llama-2-70B and Mistral-7B-v0.1 (both originally included in Lin et al. (2024), unlike most others). Conversely, the 2nd-turn score with Urial is significantly worse than for instruction-tuned LLMs: we hypothesize that this is because no multi-turn demonstrations are given in context. Indeed, Zhou et al. (2023) show that, even in IFT, having a few multi-turn training examples significantly improve the performance of the model on multi-turn conversations. Because of this fact, in the rest of the paper, we mainly focus on the 1st-turn score and single-round conversations, and track of 2nd-turn performance for completeness. To fill in the gap in the influence of multi-turn conversations, we study multi-turn alignment via ICL in Sec. 2.4. Finally, we detail the breakdown of the results over categories in Table 11 (see App. E.4 for details).
@@ 2.2 What matters for in-context alignment via Urial?
Figure 1: Effect of decoding parameters on the 1st-turn MT-Bench scores across models. We vary temperature and top-$p$, with a fixed repetition penalty of 1.15 as used in the Urial codebase. The heatmaps show that the answering quality of the base model (Mistral-7B-v0.2) with and without Urial is sensitive to the decoding schemes. Conversely, the performance of the instruct model is robust to varying the decoding parameters. Surprisingly, with proper decoding parameters, the base model alone is already capable of following instructions. See the complete results in App. E.3.
Decoding parameters. When computing the performance of Urial, Lin et al. (2024) use decoding parameters (temperature $=0$, top-$p=1$, repetition penalty $=1.15$) which differ from the default ones in MT-Bench (where temperature depends on the topic of the question, top-$p=1$, repetition penalty $=1$). To clarify the influence of these parameters that were not explicitly discussed in Lin et al. (2024), we compute the performance of base LLM, base LLM with Urial, and fine-tuned LLM when varying decoding configurations. In Fig. 1 we fix the repetition penalty to $1.15$, and vary temperature and top-$p$, for the Mistral-7B-v0.2 models. Surprisingly, we notice that with the decoding parameters of Urial (temperature $=0$, top-$p=1$) even the base model without any in-context example achieves reasonable MT-Bench score (6.61 vs 7.00 of Urial). Moreover, the temperature value appears to have the most influence of the performance of the base model. With Urial, almost all configurations provide very similar results, with the exception of temperature $=1$, top-$p=1$ (possibly because it allows the sampling of generated tokens from the tail of the token distribution, thus producing low-quality text). Finally, the aligned model performs similarly across all decoding schemes, with slightly better results than Urial. These results suggest that (1) the decoding parameters are an overlooked factor contributing to the success of in-context alignment, and (2) fine-tuning adjusts the sampling distribution of language models so that even with high-variance decoding configurations the generated text preserves high quality. We provide additional analysis for no repetition penalty (i.e. $=1$) in Fig. 13 and Llama-3.1-8B in Fig. 14 in App. E.3: these extensive experiments further validate, with different base LLMs and settings, that the decoding scheme is a crucial factor for the instruction-following behavior of base models, and also affects, but to lower degree, in-context alignment. Instead, the fine-tuned models are robust to variations in the decoding parameters. Then, for the rest of the experiments we fix the decoding scheme of Urial.
@@ 2.3 Importance of question-answer matching
2.4 Multi-turn alignment via ICL
Table 3: In-context alignment with multi-turn examples. Second-turn answers for all methods are generated using GPT-4o. Extending URIAL with two-turn examples improves the 2nd-turn MT-Bench score. This shows that in-context alignment is not limited to single-turn performance.
In-context prompt Mistral-7B-v0.2 Llama-3.1-8B
Turn 1 Turn 2 Avg Turn 1 Turn 2 Avg
Base model (no prompt) $6.52$ $5.68$ $6.10$ $6.25$ $5.17$ $5.71$
Original Urial (single-turn) $6.99$ $5.55$ $6.27$ $6.95$ $5.31$ $6.13$
Two-turn Urial $7.18$ $5.68$ $6.43$ 7.31 $6.44$ $6.87$
Urial + 3 two-turn SkillMix examples (no separation tag) 7.46 $5.88$ $6.67$ $7.21$ $6.43$ $6.82$
Urial + 3 two-turn SkillMix examples (separation tag) $7.20$ 6.31 6.76 $7.22$ 6.71 6.97
Urial + 6 two-turn SkillMix examples (no separation tag) $7.18$ $5.89$ $6.53$ $7.14$ $6.26$ $6.70$
Urial + 6 two-turn SkillMix examples (separation tag) $7.24$ $5.78$ $6.51$ $6.89$ $6.33$ $6.61$
In IFT, Zhou et al. (2023) find that adding a few multi-turn examples to the training data significantly improves the model’s performance on multi-turn conversations. Urial (Lin et al., 2024) shows that the base LLMs can handle multi-turn conversations even with only having single-turn demonstrations in the context, but it does not quantify the performance nor rigorously study the role of examples that have more than one turn of conversation. To bridge this research gap, we further test how the presence of multi-turn examples in the context influences the instruction-following performance of in-context aligned models.
Experimental setup. We experiment using two open source LLMs, Mistral-7B-v0.2 and Llama-3.1-8B, and evaluate on MT-Bench. To ensure the multi-turn examples we test have high quality, we select single-turn examples randomly from SkillMix as the basis, the first turn, and then prompt GPT-4o to ask a follow-up question after seeing the first-turn conversation and generate an answer by using an instruction ”Here is a single-turn conversation between a user and an assistant, please ask a follow-up question and provide an answer.”. We examine the influence of two-turn examples from two perspectives, the number of two-turn examples (3 vs 6), and the format, with or without separation tags like ”Here are 3 single-turn/multi-turn examples:”.
Results. In Table 3, we observe that extending 3 Urial single-turn examples to 2-turn examples consistently improves the evaluation scores on both turns, showing that in-context alignment solely with multi-turn examples can take care of single-turn conversations. The presence of two-turn examples in the context always leads to improved 2nd-turn MT-Bench scores, for example, an increase from 5.31 to 6.43 after adding 3 two-turn SkillMix examples on top of Urial on Llama-3.1-8B. Another finding from our experiments is that the format of placing demonstrations plays a pivotal role in multi-turn interactions. In particular, adding a separation tag to arrange demonstrations in two groups improves 2nd-turn scores, from 5.88 to 6.31 on Mistral-7B-v0.2 and from 6.43 to 6.71 on Llama-3.1-8B, although sometimes it hurts the 1st-turn performance. However, we find that increasing the number of two-turn examples in the context does not lead to more benefits with or without a better format.
@@ 3 Many-Shot In-Context Learning for Instruction Following
Data. We adopt the SkillMix (Kaur et al., 2024) dataset consisting 4,000 examples as the source of high-quality examples for our many-shot in-context alignment experiments. The query-answer paired data generation process of SkillMix explicitly combines diversity and difficulty through random skill combinations, where the set of core skills for instruction-following, such as developing training programs and digital proficiency, are extracted by prompting a powerful LLM like GPT-4-Turbo. We present more details of SkillMix, including some examples (Fig. 6 and Fig. 7), in App. A.1.
@@ 3.1 Scaling up the number of in-context demonstrations
Finally, we notice that on Llama-3.1-8B, using three examples from SkillMix attain, on average, better performance than using the three examples of Urial. Overall, there is no significant difference in the scaling behavior between using or not using the Urial demonstrations. Thus, we conclude that the success of in-context alignment is only partially dependent on the demonstrations themselves, provided they are of sufficient quality (see also Sec. 4). We also include the results of scaling experiments (Sec. E.1) conducted on AlpacaEval 2.0, which consists of 805 examples.
@@ 3.2 Greedy search for effective demonstrations
We add the results of the greedy search to the plots in Fig. 3, computing the true MT-Bench (i.e., with GPT-4 as judge). For both base models, the 4th example found by greedy search is sufficient to match the best 1st-turn score achieved by scaling the in-context examples with random demonstrations (see Sec. 3.1). For Llama-3.1-8B, the 5th and 6th demonstrations further improve the score, while they are not helpful for Mistral-7B-v0.2. In Table 4, we provide the details of these results: the 4th ICL example yields a significant improvement over Urial, e.g., from 6.95 to 7.61 for Llama-3.1-8B, with only a further +0.20 given by the 5th and 6th. Overall, this evaluation further indicates that the improvement in instruction following one can achieve with in-context alignment quickly saturates when increasing the number of examples. In Fig. 4, we show the distribution of the scores (with GPT-4-Turbo as judge) at step of the greedy search when using Llama-3.1-8B as base model (the analog for Mistral-7B-v0.2 in App. E.2). Most demonstrations improve the score when added as the 4th example on top of Urial (as shown in the first plot), but no further progress is observed with additional demonstrations.
@@ 5 Conclusions
In this work, we have first illustrated that ICL via Urial is a good baseline for instruction-following alignment, but with a few limitations: it typically performs slightly worse than IFT on single-turn conversations and does not generalize well to multi-round ones. Then, we have uncovered the key components for IC alignment: e.g. the decoding parameters alone can, surprisingly, make base models reasonably good at instruction following. Adding in-context high-quality demonstrations improves performance beyond what previously suggested (Lin et al., 2024), but not enough to reach LLMs aligned with sophisticated methods (e.g., RLHF). We conjecture that via ICL, the LLM can learn to infer the response style, but the overall learning signal is not sufficiently strong to benefit from a large amount of examples, despite the long context windows of recent LLMs.
@@ Acknowledgements
We are grateful to OpenAI for providing us API credits and access to GPT-4-Base as part of the OpenAI Researcher Access Program. M.A. was supported by the Google Fellowship and Open Phil AI Fellowship. This work was partially funded by an unrestricted gift from Google and by the Swiss National Science Foundation (grant number 212111).
@@ Appendix A Experimental Details
Figure 6: An example of SkillMix describes a conversation about planning and includes two skills, developing training programs and digital proficiency.
Figure 7: An example of SkillMix describes a conversation about interpretive explanation and includes two skills, performance metrics analysis and writing review articles.
@@ A.1 Datasets and base models
We select instruction-following demonstrations for the base LLMs to learn in context from open-sourced instruction fine-tuning datasets: (a) SkillMix-4k (Kaur et al., 2024) is an automated approach for synthesizing diverse, high-quality IFT data. It primarily involves two stages: first leveraging the LLMs to propose a set of critical ”skills” for instruction-following, from which a pair of skills are randomly chosen to facilitate synthetic data generation based on powerful LLMs. We show two examples from SkillMix in Fig. 6 and Fig. 7. (b) Evol-Instruct-70k (Xu et al., 2024) contains 70k training examples with varying complexity and is well-known for the use to build the series of WizardLM models. SkillMix-4k is the primary dataset for all experiments in the paper and Evol-Instruct-70k is solely used for scaling experiments and ablation study of data quality.